Check the dimensional correctness of the following
Pressure P = ( ρgh )1/2
Where ρ is the density, g is the acceleration due to gravity, h is the height
1
Expert's answer
2020-11-23T10:27:58-0500

Solution

According to given data

Pressure P = ( ρgh )1/2

Left hand side

"pressure=\frac{N}{m^2}=
\frac{Kg - m}{m^2- sec^2}\\=\frac{Kg }{m-sec^2}"

Now Right hand side

"\rho gh=\frac{Kg }{m^3}.\frac{m }{sec^2}.m"

="\frac{Kg }{m- sec^2}"

So left and right hand side are equal in dimensional.